OpenStudy (anonymous):

Find a and b. y= ax^2 + bx - 27; (2, -3) I can't seem to find b that's my problem...

jimthompson5910 (jim_thompson5910):

is (2,-3) the vertex?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

do you know vertex form?

OpenStudy (anonymous):

If I can find what be is then I'm all set. Yeah, ax^2 to bx + C

OpenStudy (anonymous):

To find be would be -b/2a =2?

OpenStudy (anonymous):

I can only get to -3=a(2)^2+b(2)+C

jimthompson5910 (jim_thompson5910):

vertex form is y = a(x-h)^2 + k we know the vertex is (2,-3), so h = 2 and k = -3 This means we get y = a(x - 2)^2 - 3 We also know that the y-intercept is (0, -27) after plugging in x = 0 into the original equation and getting y = -27, so plug in x = 0 and y = -27 to get -27 = a(0 - 2)^2 - 3 Solve for 'a' to get -27 = a(0 - 2)^2 - 3 -27 = a(-2)^2 - 3 -27 = a(4) - 3 -27 = 4a - 3 4a - 3 = -27 4a = -27+3 4a = -24 a = -24/4 a = -6 ------------------------------- So we now know that y = a(x - 2)^2 - 3 becomes y = -6(x - 2)^2 - 3 To find 'b', expand this last equation out and compare it to y= ax^2 + bx - 27

OpenStudy (anonymous):

Oh my gosh. Thank you so much! I never thought of using that equation. =) So helpful.

jimthompson5910 (jim_thompson5910):

you're welcome